OpenStudy (anonymous):

help please !!!!!!!!!!! 7x 2x --- - --- 3x^2-2x-16 x^2+x-2

OpenStudy (anonymous):

find the difference

OpenStudy (anonymous):

is --- that division?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

you can start out by factoring the bottoms

OpenStudy (anonymous):

3x^2-2x-16 = (x+2)(3x-8)

OpenStudy (anonymous):

x^2+x-2= (x-1)(x+2)

OpenStudy (anonymous):

so what would be the numerator ? the bottom is (3x-8)(x-1)

OpenStudy (anonymous):

okay sorry i wish the tools were workiing so i could show you

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%287x%29%2F%283x%5E2-2x-16%29-%282x%29%2F%28x%5E2%2Bx-2%29 there is the correct answer.

OpenStudy (anonymous):

thank you soooo much. That wedsite was very helpfull. thank you

OpenStudy (anonymous):

no problem, any time!